# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 3.2.1 CNN-based RSISR
Others CNN-based Methods. In addition to the above-mentioned CNN-based methods, some other CNN-based RSISR approaches have also been proposed. For example, Chang and Luo [190] proposed a bidirectional ConvLSTM network, which reuses local low-level features through a recursive inference block and combines a bidirectional ConvLSTM layer to adaptively select complementary information. Zhu et al. [191] proposed a degradation model based on satellite imaging and ground post-processing to improve the SR performance of real satellite images. Deeba et al. [192] proposed the Wide Remote Sensing Residual Network (WRSR), which optimizes the efficiency of single-image super-resolution by increasing the width of the residual network, reducing the depth, and using weight normalization to reduce memory costs. Wei and Liu [193] introduced dilated convolution into residual dense blocks while keeping parameters and receptive fields unchanged, and combined them with cloud services to optimize remote sensing image reconstruction. In addition, Zhao et al. [194] utilized super-Laplacian gradient prior to constrain SR reconstruction and enhanced geometric structure recovery through second-order supervision. Some recent studies have used multimodal fusion to improve performance. For example, the DeepOSWSRM method proposed by Yin et al. [195] fuses Sentinel-1 SAR and Sentinel-2 optical images to solve the cloud occlusion problem and further generate high-resolution water distribution maps. In addition, Dai et al. [196] utilized the polarization information and spatial features of polarimetric SAR to extract and fuse polarization context features through a dual-branch architecture to improve the super-resolution performance of PolSAR images.
@@ 3.2.2 GAN-based RSISR
In 2017, Ledig et al. [197] first applied generative adversarial networks (GANs) to image super-resolution. Since then, many GAN-based RSISR methods have been proposed, and GAN-based methods have become one of the most important types of RSISR methods. In general, The GAN-based RSISR methods consist of generators and discriminators. The generator reconstructs the LR image into an HR image, and the discriminator determines whether the image is a true HR (GT HR). Specifically, a two-stage training strategy is adopted in this type of methods. First, the generator is frozen to train the discriminator, thereby enhancing its discrimination ability. Then the discriminator is frozen to train the generator, and the generation result is continuously optimized through adversarial feedback. At the same time, reconstruction loss and perceptual loss are used to improve image structure restoration and visual quality.
@@
Network architecture optimization. Some studies improve the generator and discriminator structures to enhance feature extraction and reconstruction capabilities. For example, Ma et al. [198] reduced the computational burden by removing the batch normalization layer, and used transfer learning to alleviate the problem of insufficient data. Subsequently, in their follow-up research, the DRGAN method [199] designed a dense residual generator and combined it with Wasserstein GAN to improve training stability. Guo et al. [70] proposed a cascaded GAN method for RSISR, which introduced scene constraints and edge enhancement modules to improve performance. In addition, Wang et al. [200] innovatively constructed a two-dimensional matrix topology structure to significantly increase the connection density of the network for RSISR. The EEGAN method [58] proposed a dual-branch network structure to suppress noise through an edge enhancement subnet. Furthermore, the TDRRDB-EEGAN method [201] improved the edge purification module of EEGAN and combined the intermediate image with the enhanced edge to generate clear results. Pang et al. [202] replaced the generator with ResNet-50 and inserted a fully connected layer in the discriminator, significantly improving the quality of super-resolution images while maintaining the reconstruction speed. Liu et al. [203] proposed a cascaded conditional Wasserstein GAN combined with residual dense blocks to optimize gradient propagation. Guo et al.[204] designed a multi-level dense network and matrix mean discriminator to optimize the reconstruction of real aerial images. Similarly, Sustika et al. [205] combined the residual dense network (RDN) with GAN to balance objective indicators and perceptual quality. In 2023, the MSRBGAN method proposed by Zhao et al. [206] introduced multi-scale residual blocks and GAM attention to improve the performance of RSISR. The DRGAN method in recent work [112] proposed dynamic convolution and self-attention mechanism to further optimize feature extraction.
@@
Attention Mechanisms. Attention mechanisms and multi-scale modeling are used in some studies. A typical example is that SCSE-GAN [207] first introduced spatial-channel attention into the generator to enhance feature representation. In addition, Li et al. [208] improved performance by capturing long-range dependencies based on local and global attention. Zhang et al.[209] proposed a coordinated attention mechanism and residual stacking strategy to optimize landslide scene reconstruction. Jia et al. [26] integrates pyramid convolution residuals, pixel attention, and branch attention to improve the performance of RSISR. Furthermore, Wang et al. [210] optimizes RSISR algorithm by combining multi-scale structure with channel-spatial attention. The CEEGAN proposed by Ren et al. [211] designs an edge feature enhancement module (EFEM) to fuse multi-scale contextual information and optimize edge recovery for RSISR. In 2024, MFFAGAN proposed by Tang et al. [212] developed a multi-level feature fusion module to enhance key information extraction through hybrid attention. The TDEGAN method [213] introduces Shuffle attention and artifact loss functions to effectively restore high-frequency textures. Similarly, SRDSRAN [99] and SRGAN-MSAM-DRC [214] use layered dense sampling and multi-scale attention mechanisms to significantly improve the performance of large-scale RSISR.
@@
Loss Functions and Training Strategies. Some studies have focused on optimizing loss functions and training strategies to further improve the performance of RSISR. In 2019, Lei et al. [67] proposed a coupled discriminator to alleviate the low-frequency region discrimination ambiguity through a dual-path network and a random gate mechanism. In addition, Fan et al. [215] used a U-Net discriminator combined with a CBAM module to enhance detail discrimination for RSISR. Similarly, a RGTGAN method proposed by Tu et al. [27] designed a gradient-assisted texture enhancement module (GTEM) and a dense-intern deformable convolution (DIDConv) to improve cross-view alignment for RSISR. On the other hand, the RWDM method [77] constructs a real degradation training dataset and combines it with the residual balanced attention network (RBAN) to improve the reconstruction effect of the actual scene. In addition, the BLG-GAN method [216] uses a two-stage training strategy to migrate real LR images to the bicubic degradation domain to alleviate the simulation data bias. MCWESRGAN method [217] uses Wasserstein loss and multi-column discriminator to accelerate training for RSISR. Zhu et al. [218] introduced the relative average GAN (RaGAN) reconstruction loss to optimize edge detail learning, and Guo et al. [219] improved texture fidelity based on Charbonnier loss and semantic perception loss.
@@
Cross-modal and Multi-task fusion. Recently, cross-modal methods have been studied. For example, Kong et al. [220] combined Planet Fusion with Landsat data to generate high spatiotemporal resolution NDVI series. In addition, Sun et al. [221] fused high-resolution optical images with digital surface models (DSM) to preserve terrain features through slope loss. The VHR-GAN method [222] uses PeruSat1 satellite data to achieve super-resolution of Sentinel-2 images. In addition, Huang Hejing [223] combines wavelet transform to optimize frequency domain feature expression. The MRSISR method [224] proposes a multi-frame super-resolution framework that uses temporal information to optimize video satellite image reconstruction. Some other examples are the Enlighten-GAN method [225] and recent work [226] which both integrate spatial image modality and frequency domain modality information, and optimize the RSISR algorithm by combining perceptual loss with pixel reconstruction loss.
@@ 3.2.3 Transformer-based RSISR
In the field of RSISR, the conventional CNN-based methods still face challenges due to the inherent properties of remote sensing images, such as the coexistence of multi-scale targets, complex spectral-spatial coupling characteristics, and low processing efficiency of large-size images. In addition, the conventional CNN-based approaches often struggle to capture global contextual information due to their limited local receptive fields, frequently resulting in edge blurring or undesirable artifacts. In contrast, Transformer architecture’s global attention mechanism offers a promising solution to address these limitations. To the best of our knowledge, Ye et al. [229] proposed the first Transformer-based RSISR method, which combines CNN and Transformer modules to achieve efficient super-resolution reconstruction of LR remote sensing images. Since then, various Transformer-based RSISR methods have been proposed, including transform-domain-based, multi-scale features, multi-modal fusion, special modality input, and improved self-attention mechanisms, and the detailed survey in this section is as follows:
@@
Multi-Scale Features. Many studies considered the complexity of remote sensing image scenes and utilized multi-scale feature fusion to improve performance. Specifically, multi-scale feature fusion methods effectively balance the global structure coherence and local detail authenticity through dynamic weight allocation. Early studies such as TransENet proposed by Lei et al. [79] built a multi-stage enhanced structure based on Transformer, which embeds and fuses multi-scale high/low-dimensional features through an encoder-decoder architecture, breaking through the limitation of traditional models that rely on upsampling layers and ignore high-dimensional spatial features. In addition, the MSFNet proposed by Shi et al. [230] achieved hierarchical feature propagation and adaptive cross-scale fusion across iteration stages based on multi-scale implicit constraints and row-column decoupling Transformer modules. Shang et al. [231] proposed a hybrid-scale hierarchical Transformer network (HSTNet) for RSISR, which explicitly modeled single-scale and cross-scale long-range dependencies via hybrid-scale recursive feature mining and enhanced the discriminative ability of high-dimensional features. In 2024, some studies have made breakthroughs in balancing computational efficiency and multi-scale perception. For example, SymSwin proposed by Jiao et al. [101] involved a symmetric multi-scale window mechanism (SyMW) and cross-receptive field adaptive attention (CRAA), which dynamically fuses contextual features of different sizes and enhances supervision through frequency domain wavelet transform loss (UWT). In addition, Lu et al. [232] proposed a Multi-Scale and Global Representation Enhancement-based Transformer (MSGFormer), which designed the dual-window self-attention (DWSA) and multi-scale deep convolutional attention (MDCA), and tracking back structure (TBS) to capture local-global multi-scale features at a lower computational cost. For geographic data, Wang et al. [102] introduced local-global deformable blocks (LGDB) to fuse the multi-scale heterogeneity of terrain, supplemented by spatial-channel coupled attention, which significantly alleviated the local smoothing problem in DEM reconstruction.
@@
Transform Domain. The transform domain-based methods mainly separate the high- and low-frequency information of the image via frequency domain decomposition (such as wavelet transform and Fourier transform), and utilize the global attention mechanism of Transformer to strengthen the feature correlation. For example, the CFMB-T method proposed by Cao et al. [233] in 2023 employs SwinIR as the architecture, which separated the frequency domain features of infrared remote sensing images through wavelet decomposition, and adds degradation priors and recursive modules to improve performance. The dual SR framework [234] in 2024 further combined self-supervised learning with wavelet fusion, and modeled the long-range spatiotemporal dependency of remote sensing images through the Transformer attention mechanism, effectively solving the problem of periodic resolution degradation caused by the rotating synthetic aperture system. Recently, Xie et al. [235] proposed a RSISR framework based on Swin Transformer, named CFFormer, which enhances feature expression through frequency domain channel Fourier blocks (CFB) and global attention modules (GAB), and optimizes gradients by combining jump joint fusion mechanism. In addition, Liu and Yang [236] proposed a frequency domain feature hierarchical architecture based on discrete wavelet transform (DWT), which learns the correlation between full-band and high-frequency features through a self-attention mechanism. In summary, Transform-domain-based methods usually outperform pure spatial domain models, especially showing stronger detail reconstruction capabilities in complex degradation scenarios.
@@
Special Modality-guided. Recent studies have employed special modality input to improve the performance of Transformer in RSISR tasks. For example, Cai and Zhang [238] proposed a texture transfer transformer-based RSISR (T3SR) method, which first introduced the texture transfer paradigm into the remote sensing field, and solved the single image texture loss problem through a two-stage framework of shallow texture transfer and U-Transformer feature fusion, while greatly reducing the model’s dependence on reference images. In addition, the hybrid attention model proposed by Alireza and Mahdi [109] combines multi-head attention and spatial-channel attention mechanisms, which is optimized for the multispectral characteristics of Sentinel-2, and achieves SOTA performance. GeoSR framework for drone images proposed by Zhao and Li [103] pioneered a task-driven evaluation paradigm, which uses semantic segmentation accuracy as a supplementary modality to traditional PSNR/SSIM, and establishes a positive correlation between super-resolution quality and 8.6% accuracy of object classification. These studies further illustrated the potential of special modality-guided reconstruction such as texture/gradient/semantic maps in improving RSISR performance.
@@
Improved Self-attention Mechanisms. In order to solve the problems of high computational complexity, insufficient local receptive field, and inefficient cross-stage feature fusion in the traditional global self-attention mechanism, some research based on improving the self-attention mechanism has been carried out to improve performance. Specifically, most of these studies are mainly based on enhancing context modeling capabilities, optimizing feature fusion efficiency, and reducing computational complexity. For example, Lu et al. [239] proposed a cross-spatial pixel integration and cross-stage feature fusion-based transformer network (SPIFFNet) for RSISR. Specifically, the spatial pixel integration attention (CSPIA) module is designed to introduce contextual information into the local window, and the cross-stage feature fusion attention (CSFFA) module is designed to achieve adaptive feature transfer. The ESSAformer method proposed by Zhang et al. [85] developed a kernelized self-attention (ESSA) based on spectral correlation coefficient, which reduces the computational complexity to linear and strengthens the interaction of spectral features. In addition, Mao et al. [240] proposed a distance-enhanced strip attention transformer (DESAT), which enhances spatial correlation by fusing strip window attention with distance prior, and designed an attention-enhanced upsampling module. Kang et al. [241] recently proposed an Aggregation Connection Transformer (ACT-SR) for RSISR, which constructs an aggregate connection attention block and uses a series-parallel hybrid connection to fuse spatial-channel features, supplemented by a gated feedforward network to enhance nonlinear expression capabilities. In addition, the scale-aware backprojection Transformer (SPT) proposed by Hao et al. [242] innovatively combined backprojection learning with Transformer to construct a scale-aware self-attention layer, thereby improving feature learning efficiency. Kang et al. [28] proposed an efficient Swin Transformer (ESTNet) via channel attention, which reduces the number of parameters by 83% by synergizing group attention with efficient channel attention. In recent work, Xiao et al. [29, 30] proposed a Top-k Token Selective Transformer (TTST) for RSISR, which designed a three-level self-attention optimization framework involving Residual Token Selective Group (RTSG), multi-scale feature fusion (MFL) and Global Context Attention (GCA), further solving the redundant calculation and single-scale modeling limitations of self-attention in specific tasks.

### v3

@@ 3.2.1 CNN-based RSISR
Others CNN-based Methods. In addition to the above-mentioned CNN-based methods, some other CNN-based RSISR approaches have also been proposed. For example, Chang and Luo [193] proposed a bidirectional ConvLSTM network, which reuses local low-level features through a recursive inference block and combines a bidirectional ConvLSTM layer to adaptively select complementary information. Zhu et al. [194] proposed a degradation model based on satellite imaging and ground post-processing to improve the SR performance of real satellite images. Deeba et al. [195] proposed the Wide Remote Sensing Residual Network (WRSR), which optimizes the efficiency of single-image super-resolution by increasing the width of the residual network, reducing the depth, and using weight normalization to reduce memory costs. Wei and Liu [196] introduced dilated convolution into residual dense blocks while keeping parameters and receptive fields unchanged, and combined them with cloud services to optimize RSI reconstruction. In addition, Zhao et al. [197] utilized super-Laplacian gradient prior to constrain SR reconstruction and enhanced geometric structure recovery through second-order supervision. Some recent studies have used multimodal fusion to improve performance. For example, the DeepOSWSRM method proposed by Yin et al. [198] fuses Sentinel-1 SAR and Sentinel-2 optical images to solve the cloud occlusion problem and further generate high-resolution water distribution maps. In addition, Dai et al. [199] utilized the polarization information and spatial features of polarimetric SAR to extract and fuse polarization context features through a dual-branch architecture to improve the super-resolution performance of PolSAR images.
@@ 3.2.2 GAN-based RSISR
In 2017, Ledig et al. [200] first applied generative adversarial networks (GANs) to image super-resolution. Since then, many GAN-based RSISR methods have been proposed, and GAN-based methods have become one of the most important types of RSISR methods. In general, the GAN-based RSISR methods consist of generators and discriminators. The generator reconstructs the LR image into an HR image, and the discriminator determines whether the image is a true HR (GT HR), as shown in Fig. 8. Specifically, a two-stage training strategy is adopted in this type of method. First, the generator is frozen to train the discriminator, thereby enhancing its discrimination ability. Then the discriminator is frozen to train the generator, and the generation result is continuously optimized through adversarial feedback. At the same time, reconstruction loss and perceptual loss are used to improve image structure restoration and visual quality.
@@
Network architecture optimization. Some studies improve the generator and discriminator structures to enhance feature extraction and reconstruction capabilities. For example, Ma et al. [201] reduced the computational burden by removing the batch normalization layer and used transfer learning to alleviate the problem of insufficient data. Subsequently, in their follow-up research, the DRGAN method [202] designed a dense residual generator and combined it with Wasserstein GAN to improve training stability. Guo et al. [71] proposed a cascaded GAN method for RSISR, which introduced scene constraints and edge enhancement modules to improve performance. In addition, Wang et al. [203] innovatively constructed a two-dimensional matrix topology structure to significantly increase the connection density of the network for RSISR. The EEGAN method [59] proposed a dual-branch network structure to suppress noise through an edge enhancement subnet. Furthermore, the TDRRDB-EEGAN method [204] improved the edge purification module of EEGAN and combined the intermediate image with the enhanced edge to generate clear results. Pang et al. [205] replaced the generator with ResNet-50 and inserted a fully connected layer in the discriminator, significantly improving the quality of super-resolution images while maintaining the reconstruction speed. Liu et al. [206] proposed a cascaded conditional Wasserstein GAN combined with residual dense blocks to optimize gradient propagation. Guo et al.[207] designed a multi-level dense network and matrix mean discriminator to optimize the reconstruction of real aerial images. Similarly, Sustika et al. [208] combined the residual dense network (RDN) with GAN to balance objective indicators and perceptual quality. In 2023, the MSRBGAN method proposed by Zhao et al. [209] introduced multi-scale residual blocks and GAM attention to improve the performance of RSISR. The DRGAN method in recent work [113] proposed a dynamic convolution and a self-attention mechanism to further optimize feature extraction.
@@
Attention Mechanisms. Attention mechanisms and multi-scale modeling are used in some studies. A typical example is that SCSE-GAN [210] first introduced spatial-channel attention into the generator to enhance feature representation. In addition, Li et al. [211] improved performance by capturing long-range dependencies based on local and global attention. Zhang et al.[212] proposed a coordinated attention mechanism and residual stacking strategy to optimize landslide scene reconstruction. Jia et al. [26] integrate pyramid convolution residuals, pixel attention, and branch attention to improve the performance of RSISR. Furthermore, Wang et al. [213] optimizes RSISR algorithm by combining multi-scale structure with channel-spatial attention. The CEEGAN proposed by Ren et al. [214] designs an edge feature enhancement module (EFEM) to fuse multi-scale contextual information and optimize edge recovery for RSISR. In 2024, MFFAGAN proposed by Tang et al. [215] developed a multi-level feature fusion module to enhance key information extraction through hybrid attention. The TDEGAN method [216] introduces Shuffle attention and artifact loss functions to effectively restore high-frequency textures. Similarly, SRDSRAN [100] and SRGAN-MSAM-DRC [217] use layered dense sampling and multi-scale attention mechanisms to significantly improve the performance of large-scale RSISR.
@@
Loss Functions and Training Strategies. Some studies have focused on optimizing loss functions and training strategies to further improve the performance of RSISR. In 2019, Lei et al. [68] proposed a coupled discriminator to alleviate the low-frequency region discrimination ambiguity through a dual-path network and a random gate mechanism. In addition, Fan et al. [218] used a U-Net discriminator combined with a CBAM module to enhance detail discrimination for RSISR. Similarly, a RGTGAN method proposed by Tu et al. [27] designed a gradient-assisted texture enhancement module (GTEM) and a dense-intern deformable convolution (DIDConv) to improve cross-view alignment for RSISR. On the other hand, the RWDM method [78] constructs a real degradation training dataset and combines it with the residual balanced attention network (RBAN) to improve the reconstruction effect of the actual scene. In addition, the BLG-GAN method [219] uses a two-stage training strategy to migrate real LR images to the bicubic degradation domain to alleviate the simulation data bias. MCWESRGAN method [220] uses Wasserstein loss and a multi-column discriminator to accelerate training for RSISR. Zhu et al. [221] introduced the relative average GAN (RaGAN) reconstruction loss to optimize edge detail learning, and Guo et al. [222] improved texture fidelity based on Charbonnier loss and semantic perception loss.
@@
Cross-modal and Multi-task fusion. Recently, cross-modal methods have been studied. For example, Kong et al. [223] combined Planet Fusion with Landsat data to generate high spatiotemporal resolution NDVI series. In addition, Sun et al. [224] fused high-resolution optical images with digital surface models (DSM) to preserve terrain features through slope loss. The VHR-GAN method [225] uses PeruSat1 satellite data to achieve super-resolution of Sentinel-2 images. In addition, Huang Hejing [226] combines the wavelet transform to optimize the frequency domain feature expression. The MRSISR method [227] proposes a multi-frame super-resolution framework that uses temporal information to optimize video satellite image reconstruction. Some other examples are the Enlighten-GAN method [228] and recent work [229], which both integrate spatial image modality and frequency domain modality information, and optimize the RSISR algorithm by combining perceptual loss with pixel reconstruction loss.
@@ 3.2.3 Transformer-based RSISR
In the field of RSISR, the conventional CNN-based methods still face challenges due to the inherent properties of RSIs, such as the coexistence of multi-scale targets, complex spectral-spatial coupling characteristics, and low processing efficiency of large-size images. In addition, the conventional CNN-based approaches often struggle to capture global contextual information due to their limited local receptive fields, frequently resulting in edge blurring or undesirable artifacts. In contrast, Transformer architecture’s global attention mechanism offers a promising solution to address these limitations. To the best of our knowledge, Ye et al. [232] proposed the first Transformer-based RSISR method, which combines CNN and Transformer modules to achieve efficient super-resolution reconstruction of LR RSIs. Since then, various Transformer-based RSISR methods have been proposed, including transform-domain-based, multi-scale features, multi-modal fusion, special modality input, improved self-attention mechanisms, and the detailed survey in this section is as follows:
@@
Multi-Scale Features. Many studies considered the complexity of RSI scenes and utilized multi-scale feature fusion to improve performance. Specifically, multi-scale feature fusion methods effectively balance the global structure coherence and local detail authenticity through dynamic weight allocation. Early studies such as TransENet proposed by Lei et al. [80] built a multi-stage enhanced structure based on Transformer, which embeds and fuses multi-scale high/low-dimensional features through an encoder-decoder architecture, breaking through the limitation of traditional models that rely on upsampling layers and ignore high-dimensional spatial features. In addition, the MSFNet proposed by Shi et al. [233] achieved hierarchical feature propagation and adaptive cross-scale fusion across iteration stages based on multi-scale implicit constraints and row-column decoupling Transformer modules. Shang et al. [234] proposed a hybrid-scale hierarchical Transformer network (HSTNet) for RSISR, which explicitly modeled single-scale and cross-scale long-range dependencies via hybrid-scale recursive feature mining and enhanced the discriminative ability of high-dimensional features. In 2024, some studies have made breakthroughs in balancing computational efficiency and multi-scale perception. For example, SymSwin proposed by Jiao et al. [102] involved a symmetric multi-scale window mechanism (SyMW) and cross-receptive field adaptive attention (CRAA), which dynamically fuses contextual features of different sizes and enhances supervision through frequency domain wavelet transform loss (UWT). In addition, Lu et al. [235] proposed a Multi-Scale and Global Representation Enhancement-based Transformer (MSGFormer), which designed the dual-window self-attention (DWSA) and multi-scale deep convolutional attention (MDCA), and tracking back structure (TBS) to capture local-global multi-scale features at a lower computational cost. For geographic data, Wang et al. [103] introduced local-global deformable blocks (LGDB) to fuse the multi-scale heterogeneity of terrain, supplemented by spatial-channel coupled attention, which significantly alleviated the local smoothing problem in DEM reconstruction.
@@
Transform Domain. The transform domain-based methods mainly separate the high- and low-frequency information of the image via frequency domain decomposition (such as wavelet transform and Fourier transform), and utilize the global attention mechanism of Transformer to strengthen the feature correlation. For example, the CFMB-T method proposed by Cao et al. [236] in 2023 employs SwinIR as the architecture, which separates the frequency domain features of infrared RSIs through wavelet decomposition, and adds degradation priors and recursive modules to improve performance. The dual SR framework [237] in 2024 further combined self-supervised learning with wavelet fusion, and modeled the long-range spatiotemporal dependency of RSIs through the Transformer attention mechanism, effectively solving the problem of periodic resolution degradation caused by the rotating synthetic aperture system. Recently, Xie et al. [238] proposed a RSISR framework based on Swin Transformer, named CFFormer, which enhances feature expression through frequency domain channel Fourier blocks (CFB) and global attention modules (GAB), and optimizes gradients by combining a jump joint fusion mechanism. In addition, Liu and Yang [239] proposed a frequency domain feature hierarchical architecture based on discrete wavelet transform (DWT), which learns the correlation between full-band and high-frequency features through a self-attention mechanism. In summary, Transform-domain-based methods usually outperform pure spatial domain models, especially showing stronger detail reconstruction capabilities in complex degradation scenarios.
@@
Special Modality-guided. Recent studies have employed special modality input to improve the performance of Transformer in RSISR tasks. For example, Cai and Zhang [241] proposed a texture transfer transformer-based RSISR (T3SR) method, which first introduced the texture transfer paradigm into the remote sensing field, and solved the single image texture loss problem through a two-stage framework of shallow texture transfer and U-Transformer feature fusion, while greatly reducing the model’s dependence on reference images. In addition, the hybrid attention model proposed by Alireza and Mahdi [110] combines multi-head attention and spatial-channel attention mechanisms, which is optimized for the multispectral characteristics of Sentinel-2, and achieves SOTA performance. GeoSR framework for drone images proposed by Zhao and Li [104] pioneered a task-driven evaluation paradigm, which uses semantic segmentation accuracy as a supplementary modality to traditional PSNR/SSIM, and establishes a positive correlation between super-resolution quality and 8.6% accuracy of object classification. These studies further illustrated the potential of special modality-guided reconstruction, such as texture/gradient/semantic maps in improving RSISR performance.
@@
Improved Self-attention Mechanisms. In order to solve the problems of high computational complexity, insufficient local receptive field, and inefficient cross-stage feature fusion in the traditional global self-attention mechanism, some research based on improving the self-attention mechanism has been carried out to improve performance. Specifically, most of these studies are mainly based on enhancing context modeling capabilities, optimizing feature fusion efficiency, and reducing computational complexity. For example, Lu et al. [242] proposed a cross-spatial pixel integration and cross-stage feature fusion-based transformer network (SPIFFNet) for RSISR. Specifically, the spatial pixel integration attention (CSPIA) module is designed to introduce contextual information into the local window, and the cross-stage feature fusion attention (CSFFA) module is designed to achieve adaptive feature transfer. The ESSAformer method proposed by Zhang et al. [86] developed a kernelized self-attention (ESSA) based on spectral correlation coefficient, which reduces the computational complexity to linear and strengthens the interaction of spectral features. In addition, Mao et al. [243] proposed a distance-enhanced strip attention transformer (DESAT), which enhances spatial correlation by fusing strip window attention with a distance prior, and designed an attention-enhanced upsampling module. Kang et al. [244] recently proposed an Aggregation Connection Transformer (ACT-SR) for RSISR, which constructs an aggregate connection attention block and uses a series-parallel hybrid connection to fuse spatial-channel features, supplemented by a gated feedforward network to enhance nonlinear expression capabilities. In addition, the scale-aware backprojection Transformer (SPT) proposed by Hao et al. [245] innovatively combined backprojection learning with Transformer to construct a scale-aware self-attention layer, thereby improving feature learning efficiency. Kang et al. [28] proposed an efficient Swin Transformer (ESTNet) via channel attention, which reduces the number of parameters by 83% by synergizing group attention with efficient channel attention. In recent work, Xiao et al. [29, 30] proposed a Top-k Token Selective Transformer (TTST) for RSISR, which designed a three-level self-attention optimization framework involving Residual Token Selective Group (RTSG), multi-scale feature fusion (MFL) and Global Context Attention (GCA), further solving the redundant calculation and single-scale modeling limitations of self-attention in specific tasks.